OpenStudy (anonymous):

What is the final balance for the investment? $2500 for 3 years at 6% compounded annually A. $3211.46 B. $2977.54 C. $2950.00 D. $2650.00

OpenStudy (anonymous):

@laibashahlakhan

OpenStudy (anonymous):

Well.. I'm bad at this money questions.. @John_ES can you please help?

OpenStudy (anonymous):

Substitute in this... you'll get it :D

OpenStudy (john_es):

sriramkumar has given the right equation, \[25000\cdot\left(1+\frac{6}{100}\right)^3=2977.54\]